Solve for x: 1/2x+3/8=3/4

Answer by checkley71(8403)

NORMALLY 1/2X IS X/2 NOT 1/(2X).
IF I'M WRONG THEN YOU'RE RIGHT.
IF I'M RIGHT THEN YOU'RE WRONG.
TAKE YOUR PICK.
X/2+3/8=3/4
X/2=3/4-3/8
X/2=(2*3-3)/8
X/2=(6-3)/8
X/2=3/8 CROSS MULTIPLY
8X=6
X=6/8=3/4 ANSWER.